# Periodic double tilings of the plane

### v2

@@ 1.1. Statement of the main results
Figure 1. The three possible configurations of isoperimetric $2$-tilings of the plane relative to a given periodic lattice.
A tiling in $\mathbb{R}^{d}$ is a collection ${\mathbf{T}}$ of measurable sets so that $\left\lvert E\cap F\right\rvert=0$ for all $E,F\in{\mathbf{T}}$, $E\neq F$, and $\left\lvert\mathbb{R}^{d}\setminus\bigcup_{E\in{\mathbf{T}}}E\right\rvert=0$. Given $N\in\mathbb{N}$, a periodic $N$-tiling is a tiling $\mathbf{T}$ such that there is a lattice $G$ and there are $N$ measurable sets $E_{1},...,E_{N}$ called generators so that (up to negligible sets), we have
@@
$\operatorname{Per}(\mathbf{T})\coloneqq\frac{1}{2}\sum_{i=k}^{N}\operatorname{Per}(E_{k}),$
where $\operatorname{Per}(E)$ is the classical Caccioppoli perimeter of a measurable set (see [1, 21]). Notice that $\operatorname{Per}(\mathbf{T})$ corresponds to the $d-1$ dimensional measure of all the mutual interfaces between the generators. Given a lattice $G$, a periodic $N$-tiling $\mathbf{T}=\{E_{k}+g\colon k=1,\dots,N,g\in G\}$ is said to be isoperimetric relative to $G$, provided that for any other periodic $N$-tiling $\mathbf{T}^{\prime}=\{E^{\prime}_{k}+g\colon k=1,\dots,N,g\in G\}$ with $\left\lvert E_{k}\right\rvert=\left\lvert E_{k}^{\prime}\right\rvert$, $k=1,\dots N$ one has
@@
    $p>0$, $E_{1}$ is a Reauleaux triangle with three arcs of curvature $p>0$, $E_{2}$ is a curvilinear ennagon which is a chipped hexagon with three flat sides and three concave edges with curvature $-p$;
@@
    $p>0$, $E_{1}$ is a strictly convex curvilinear quadrangle with the symmetries of a rectangle, the four arcs have curvature $p$, $E_{2}$ is a curvilinear octagon which is a chipped parallelogram with four flat sides and four concave edges with curvature $-p$;
@@
In particular, we have that $[0,1]\ni x\mapsto\left(\mathcal{I}(x)-\mathcal{I}(0)\right)^{2}$ is concave.
@@
Figure 2. Plot of the isoperimetric profile $\mathcal{I}(x)$ where $x\in[0,\frac{1}{2}]$ is the area of the smaller tile and $1-x$ is the area of the larger tile. On the $y$-axis, we have $y_{0}\coloneqq\mathcal{I}(0)=\sqrt[4]{12}$, and $y_{i}\coloneqq\mathcal{I}(x_{i})$ for $i=1,2$ (see Theorem 1.2). The dashed lines represent stationary but not optimal tilings. The shape of the optimal tiles is depicted for the values $x=0$ (single hexagon), $x=x_{1}$ (configurations $(3;9)$ and $(4;8)$, $x=x_{2}$ (configurations $(4;8)$ and $(6;6)$) and $x=0.5$ (configuration $(6;6)$). The transition between the three different configurations occurs at $x=x_{1}$ and $x=x_{2}$. These are the only values where we have two different isoperimetric tilings with the same areas.
Finally, we point out that it is also implied by the work [29] of the first and second named authors that two adjacent hexagons with almost the same area are isoperimetric $2$-tilings, see [29, Theorem 5.3]. However, this was achieved by a qualitative argument and the explicit value of the transition point $x_{2}$ could not be deduced (even though more than $2$-cells are allowed).
@@ 2. Classification of planar isoperimetric tilings
Apart from property (4) these results are presented, with a different notation, in [29, Theorem 5.2]. Indeed, properties (1) up to (3) have been derived for partitions obtained as minimizers of the following variational problem:
@@
Even though this was not stated, the regularity results proved in [29, Theorem 5.2] actually extend to the variational problem where the lattice $G$ is fixed:
@@
Indeed, [29, Theorem 5.2] relied on [29, Lemma 4.2] (which clearly can be extended to minimizers of (2.1), see [29, Theorem 3.2]) and also on [29, Proposition 5.1]. The latter can be readily checked to hold for minimizers of (2.1) by observing, as done for [29, Proposition 5.1], that a minimizer $D,(E_{k})$ exists satisfying
@@
  3. (iii)
    Let $v$ be a vertex of the tiling. Then there exist $C_{1}$ a connected component of $E_{1}$, $C_{2}$ a connected component of $E_{2}$ and $g_{1},g_{2},g_{3}\in G$ such that the three components having vertex $v$ are either $C_{1}+g_{1},C_{2}+g_{2},C_{1}+g_{3}$ (vertex of type 1-1-2) or $C_{1}+g_{1},C_{2}+g_{2},C_{2}+g_{3}$ (vertex of type 1-2-2).
@@
Let $C$ be a connected component of a tile and suppose that it has a flat edge $e$. This means that the other component sharing the edge $e$ has the same pressure: it can either be a translation of the same component, or else it is a component of the translation of a different generating tile which has the same pressure. It cannot be a translation of a different component of the same tile because of Proposition 2.2. If $e$ is shared by the translation $g\in G$ of the same component $C$, it means that $e-g$ is another edge of the same component which, of course, is parallel to $e$. Otherwise, the two regions have the same pressure $p=-p$ and hence $p=0$. In the case $p=0$ all the components of all the tiles are straight line segments with angles of 120 degrees, so, all the components are hexagonal polygons with edges parallel to the edges of a common regular hexagon: hence they are composed of three pairs of opposite parallel edges.
@@
If $v$ is the vertex shared by $e$ and $f$, the other vertex of $f$ must be $v+g$ because it is shared with $e^{\prime}=e+g$. Let $D$ be the component on the other side of the curved edge $f$ of $C$. Following the edges of $D$ we notice that the third edge in the vertex $v+g$ (apart from $e^{\prime}$ and $f$) must be $f+g$ and then we arrive at $v+2g$ and repeat the process. It turns out that the boundary of $D$ never closes and hence $D$ is unbounded. This is a contradiction since $D$ has finite area.
For the third point notice that the three components meeting in a vertex $v$ are of the form $C_{1}+g_{1}$, $C_{2}+g_{2}$, $C_{3}+g_{3}$ with $g_{1},g_{2},g_{3}\in G$ and $C_{1}$, $C_{2}$, $C_{3}$ connected components of either $E_{1}$ or $E_{2}$. Without loss of generality suppose that $C_{1}$ and $C_{3}$ are connected components of the same tile, say $E_{1}$. Then, by Proposition 2.3, we know that $C_{3}=C_{1}$. If $C_{2}$ were also a connected component of $E_{1}$ than also $C_{2}=C_{1}$ and the whole plane would be covered by $C_{1}+g$ with $g\in G$. So $\left\lvert E_{1}\setminus C_{1}\right\rvert=0$ and $\left\lvert E_{2}\right\rvert=0$, against the assumption that the tiling is not degenerate. So $C_{2}$ is a connected component of $E_{2}$ and the third claim follows. ∎
@@
Step 1. We claim that $C$ cannot be surrounded only by translations of $E_{2}$. Otherwise, each of the six components of the translations of $E_{2}$ surrounding $C$ would share an edge with the previous and next one. By Proposition 2.2 this means that they are all different translations of the same component $D$ of $E_{2}$. So they can be written as $D+g_{k}$ with distinct $g_{1},\dots,g_{6}\in G$ choosen so that $D+g_{1},\dots,D+g_{6}$ are ordered in counter-clockwise order around $C$. Since two adjacent translations of $D$ share an edge, each edge of the hexagon $D$ must be a translation of the opposite edge, meaning that opposite edges of $D$ have equal lengths. The vectors representing the ordered edges of $D$ can hence be denoted by $v_{1},v_{2},v_{3},-v_{1},-v_{2},-v_{3}$. Since $C$ has the edges in common with the translations of $D$ the opposite edges occur in $C$ with the same order: $-v_{1},-v_{2},-v_{3},v_{1},v_{2},v_{3}$. This means that also $C$ is a translation of $D$: $C=D+g_{1}+g_{3}-g_{2}$. This means that $C$ is contained in a translation of $E_{2}$ which is a contradiction since $C$ was chosen to be a connected component of $E_{1}$.
Step 2. We claim that the opposite edges of $C$ have equal length. By the previous step and Proposition 2.2 we know that $C$ is adjacent to a translation $C+g$ of itself, with $g\in G$. So at least one couple of opposite edges have equal length. Let $v_{1},\dots,v_{6}$ be the ordered vectors identifying the edges of $C$ so that $v_{1}+\dots+v_{6}=0$. We know that $v_{k}$ is parallel to $v_{k+3}$ (with opposite direction) for $k\in\mathbb{Z}$ (mod 6). Suppose that $v_{6}=-v_{3}$ are the vectors corresponding to the two opposite sides with equal length while $v_{4}=-\alpha v_{1}$ and $v_{5}=-\beta v_{2}$. From $v_{1}+v_{2}-\alpha v_{1}-\beta v_{2}=0$ since $v_{1}$ and $v_{2}$ are independent we obtain $01-\alpha=1-\beta=$ 0 which means that also $v_{4}=-v_{1}$ and $v_{5}=-v_{2}$. The claim follows.
@@
Step 3. Conclusion (see Figure 4). Consider a connected component $C_{1}$ of $E_{1}$ and let $g\in G$ be the translation given in Step 1 such that $C_{1}+g$ is adjacent to $C_{1}$ itself. Let $p$ be one of the two vertices in common between $C_{1}$ and $C_{1}+g$. Then, by Proposition 2.3, we know that there is a connected component $C_{2}$ of $E_{2}$ and $g_{2}\in G$ such that $C_{2}+g_{2}$ has a vertex in $p$. Two edges of the component $C_{2}+g_{2}$ are in common with $C_{1}$ and $C_{1}+g$ and hence are represented by two vectors $-v,-w$ where $v,w$ represent consecutive edges of $C_{1}$ with $v+w=g$. This means that $C_{2}+g_{2}+g$ is adjacent to $C_{2}+g_{2}$. By repeating the reasoning with $C_{1},C_{1}+g$ replaced by $C_{2}+g_{2},C_{2}+g_{2}+g$ we find $C_{3}$, a connected component of $E_{1}$, and $g_{3}\in G$ such that $C_{3}+g_{3}$ has a common vertex with $C_{2}+g_{2}$ and $C_{2}+g_{2}+g$. Moreover, $C_{3}+g_{3}+g$ is adjacent to $C_{3}+g_{3}$. Now let $u_{1},v,w,-u_{1},-v,-w$ be the vectors representing the ordered edges of $C_{1}$, let $u_{2},v,w,-u_{2},-v,-w$ be the ordered edges of $C_{2}$, and $u_{3},v,w,-u_{3},-v,-w$ be the ordered edges of $C_{3}$. Notice that $u_{1}=t_{1}u$, $u_{2}=t_{2}u$, $u_{3}=t_{3}u$ all have the same direction $u$. The hexagons $C_{1}+kg$, $C_{2}+g_{2}+kg$, $C_{3}+g_{3}+kg$ with $k\in\mathbb{Z}$ form three parallel strips with direction $g$ whose interfaces are zig-zags following the vectors $v$ and $w$. We can swap the second and third strip to obtain a new tiling with the same areas and with a whole interface removed. This is contrary to the minimality of ${\mathbf{T}}$ unless $C_{3}=C_{1}=E_{1}$. Analogously one finds $C_{2}=E_{2}$. Hence $E_{1}$ and $E_{2}$ are both connected. Uniqueness is straightforward, concluding the proof. ∎
222222221111222222222222
@@
Hence $C$ must have at least a vertex $v$ of type 1-2-2. The flat edge in this vertex has a corresponding opposite flat edge which also must have a 1-2-2 corresponding vertex $w$. Moreover the next vertex to a 1-2-2 vertex, following the curved edge, is another 1-2-2 vertex which, itself has a corresponding opposite 1-2-2 vertex.
@@
In both configurations the flat edges of $C$ can be prolonged to obtain a flat hexagon $P$ containing $C$ such that $\left\{P+g\colon g\in G\right\}$ is a tiling of $\mathbb{R}^{2}$. This means that $E_{2}=C$. Each vertex of $P$ is shared by $P$ and by $P+g$ and $P+h$ with $g,h\in G$. If the vertex of $P$ is not a vertex of $C$ we easily deduce that there is a $P\setminus C$ is contained in a component of $E_{1}$ which is a Reauleaux triangle (i.e., a curvilinear triangle with three equal angles of 120 degrees).
@@
The configuration $(3,3;12)$ is a stationary configuration but we claim that it is not isoperimetric. In fact, it is possible to increase the size of the larger Reauleaux triangle and decrease the size of the other Reauleaux triangle so that the total area enclosed by the two triangles remains constant while the perimeter decreases. This is because the contribution to the perimeter of a Reauleaux triangle (the difference between the perimeter of the triangle and the length of the corresponding tripod) is increasing and proportional to the linear size of the triangle, while the enclosed area is proportional to the square of the linear size. So, if we have two Reauleaux triangles of respective sizes $r$ and $s$, the total volume is kept fixed if $k=r^{2}+s^{2}$ is constant while the contribution to the total perimeter is proportional to
@@
Notice that the configurations $(3;9)$ is, for a general lattice, not unique because we can choose to place the Reauleaux triangle in two different vertices of the hexagon. Notice that Theorem 2.4 combined with Theorem 2.6 directly implies our first main result Theorem 1.1.

### v3

@@ 1.1. Statement of the main results
Figure 1. The three possible configurations of isoperimetric $2$-tilings of the plane relative to a given periodic lattice. See the interactive web page: https://paolini.github.io/double-tiling/.
A tiling in $\mathbb{R}^{d}$ is a collection $\mathbf{T}$} of measurable sets so that $\left\lvert E\cap F\right\rvert=0$ for all $\in{\mathbf{T}}, EFE\neq F$, and $\left\lvert\mathbb{R}^{d}\setminus\bigcup_{E\in{\mathbf{T}}}E\right\rvert=0$. Given $N\in\mathbb{N}$, a periodic $N$-tiling is a tiling $\mathbf{T}$ such that there is a lattice $G$ and there are $N$ measurable sets $E_{1},...,E_{N}$ called generators so that (up to negligible sets), we have
@@
$\operatorname{Per}(\mathbf{T})\coloneqq\frac{1}{2}\sum_{k=1}^{N}\operatorname{Per}(E_{k}),$
where $\operatorname{Per}(E)$ is the classical Caccioppoli perimeter of a measurable set (see [1, 21]). Notice that $\operatorname{Per}(\mathbf{T})$ corresponds to the d1d-1 dimensional measure of all the mutual interfaces between the generators. Given a lattice $G$, a periodic $N$-tiling $\mathbf{T}=\{E_{k}+g\colon k=1,\dots,N,g\in G\}$ is said to be isoperimetric relative to $G$, provided that for any other periodic $N$-tiling $\mathbf{T}^{\prime}=\{E^{\prime}_{k}+g\colon k=1,\dots,N,g\in G\}$ with $\left\lvert E_{k}\right\rvert=\left\lvert E_{k}^{\prime}\right\rvert$, $k=1,\dots N$ one has
@@
    $p>0$, $E_{1}$ is a Reauleaux triangle with three arcs of curvature $p>0$, $E_{2}$ is a curvilinear ennagon which is a chipped hexagon with six flat edges and three concave edges with curvature p-p;
@@
    $p>0$, $E_{1}$ is a strictly convex curvilinear quadrangle with the symmetries of a rectangle, the four arcs have curvature $p$, $E_{2}$ is a curvilinear octagon which is a chipped parallelogram with four flat sides and four concave edges with curvature p-p;
@@
In particular, we have that $0,1]\ni x\mapsto\left(\mathcal{I}(x)-\mathcal{I}(0)\right)^{2}$ is concave.
@@
Figure 2. Plot of the isoperimetric profile $\mathcal{I}(x)$ where $x\in[0,\frac{1}{2}]$ is the area of the smaller tile and 1x1-x is the area of the larger tile. On the $y$-axis, we have $y_{0}\coloneqq\mathcal{I}(0)=\sqrt[4]{12}$, and $y_{i}\coloneqq\mathcal{I}(x_{i})$ for $i=1,2$ (see Theorem 1.2 and Theorem 3.5). The dashed lines represent stationary but not optimal tilings. The shape of the optimal tiles is depicted for the values $x=0$ (single hexagon), $x=x_{1}$ (configurations $(3;9)$ and $(4;8)$, $x=x_{2}$ (configurations $(4;8)$ and $(6;6)$) and $x=0.5$ (configuration $(6;6)$). The transition between the three different configurations occurs at $x=x_{1}$ and $x=x_{2}$. These are the only values where we have two different isoperimetric tilings with the same areas.
Finally, we point out that it is also implied by the work [30] of the first and second named authors that two adjacent hexagons with almost the same area are isoperimetric $2$-tilings, see [30, Theorem 5.3]. However, this was achieved by a qualitative argument and the explicit value of the transition point $x_{2}$ could not be deduced (even though more than $2$-cells are allowed).
@@ 2. Classification of planar isoperimetric tilings
Apart from property (4) these results are presented, with a different notation, in [30, Theorem 5.2]. Indeed, properties (1) up to (3) have been derived for partitions obtained as minimizers of the following variational problem:
@@
Even though this was not stated, the regularity results proved in [30, Theorem 5.2] actually extend to the variational problem where the lattice $G$ is fixed:
@@
Indeed, [30, Theorem 5.2] relied on [30, Lemma 4.2] (which clearly can be extended to minimizers of (2.1), see [30, Theorem 3.2]) and also on [30, Proposition 5.1]. The latter can be readily checked to hold for minimizers of (2.1) by observing, as done for [30, Proposition 5.1], that a minimizer $D,(E_{k})$ exists satisfying
@@
Let $C$ be a connected component of a tile and suppose that it has a flat edge $e$. This means that the other component sharing the edge $e$ has the same pressure: it can either be a translation of the same component, or else it is a component of the translation of a different generating tile which has the same pressure. It cannot be a translation of a different component of the same tile because of Proposition 2.2. If $e$ is shared by the translation $g\in G$ of the same component $C$, it means that ege-g is another edge of the same component which, of course, is parallel to $e$. Otherwise, the two regions have the same pressure p=pp=-p and hence $p=0$. In the case $p=0$ all the components of all the tiles are straight line segments with angles of 120 degrees, so, all the components are hexagonal polygons with edges parallel to the edges of a common regular hexagon: hence they are composed of three pairs of opposite parallel edges.
@@
If $v$ is the vertex shared by $e$ and $f$, the other vertex of $f$ must be $v+g$ because it is shared with $e^{\prime}=e+g$. Let $D$ be the component on the other side of the curved edge $f$ of $C$. Following the edges of $D$ we notice that the third edge in the vertex $v+g$ (apart from $e^{\prime}$ and $f$) must be $f+g$ and then we arrive at $v+2g$ and repeat the process. It turns out that the boundary of $D$ never closes and hence $D$ is unbounded. This is a contradiction since $D$ has finite area. ∎
@@
Step 1. We claim that $C$ cannot be surrounded only by translations of $E_{2}$. Otherwise, each of the six components of the translations of $E_{2}$ surrounding $C$ would share an edge with the previous and next one. By Proposition 2.2 this means that they are all different translations of the same component $D$ of $E_{2}$. So they can be written as $D+g_{k}$ with distinct $g_{1},\dots,g_{6}\in G$ choosen so that $D+g_{1},\dots,D+g_{6}$ are ordered in counter-clockwise order around $C$. Since two adjacent translations of $D$ share an edge, each edge of the hexagon $D$ must be a translation of the opposite edge, meaning that opposite edges of $D$ have equal lengths. The vectors representing the ordered edges of $D$ can hence be denoted by $v_{1},v_{2},v_{3},-v_{1},-v_{2},-v_{3}$. Since $C$ has the edges in common with the translations of $D$ the opposite edges occur in $C$ with the same order: $v_{1},-v_{2},-v_{3},v_{1},v_{2},v_{3}$. This means that also $C$ is a translation of $D$: $C=D+g_{1}+g_{3}-g_{2}$. This means that $C$ is contained in a translation of $E_{2}$ which is a contradiction since $C$ was chosen to be a connected component of $E_{1}$.
Step 2. We claim that the opposite edges of $C$ have equal length. By the previous step and Proposition 2.2 we know that $C$ is adjacent to a translation $C+g$ of itself, with $g\in G$. So at least one couple of opposite edges have equal length. Let $v_{1},\dots,v_{6}$ be the ordered vectors identifying the edges of $C$ so that $v_{1}+\dots+v_{6}=0$. We know that $v_{k}$ is parallel to $v_{k+3}$ (with opposite direction) for $k\in\mathbb{Z}$ (mod 6). Suppose that $v_{6}=-v_{3}$ are the vectors corresponding to the two opposite sides with equal length while $v_{4}=-\alpha v_{1}$ and $v_{5}=-\beta v_{2}$. From $v_{1}+v_{2}-\alpha v_{1}-\beta v_{2}=0$ since $v_{1}$ and $v_{2}$ are independent we obtain $1-\alpha=1-\beta=0$ which means that also $v_{4}=-v_{1}$ and $v_{5}=-v_{2}$. The claim follows.
@@
Step 3. Conclusion (see Figure 4). Consider a connected component $C_{1}$ of $E_{1}$ and let $g\in G$ be the translation given in Step 1 such that $C_{1}+g$ is adjacent to $C_{1}$ itself. Let $p$ be one of the two vertices in common between $C_{1}$ and $C_{1}+g$. We claim that there is a connected component $C_{2}$ of $E_{2}$ and $g_{2}\in G$ such that $C_{2}+g_{2}$ has a vertex in $p$. If, by contradiction, it holds that $C_{2}\subset E_{1}$, then also $C_{2}=C_{1}$ (by Proposition 2.3) and the whole plane would be covered by $C_{1}+g$ with $g\in G$. So $\left\lvert E_{1}\setminus C_{1}\right\rvert=0$ and $\left\lvert E_{2}\right\rvert=0$, against the assumption that the tiling is not degenerate. So $C_{2}$ must be a connected component of $E_{2}$. Two edges of the component $C_{2}+g_{2}$ are in common with $C_{1}$ and $C_{1}+g$ and hence are represented by two vectors v,w-v,-w where $v,w$ represent consecutive edges of $C_{1}$ with v+w=gv+w=-g (recall that all connected components are hexagons with equal angles of 120 degrees and opposite edges with equal length). This means that $C_{2}+g_{2}+g$ is adjacent to $C_{2}+g_{2}$. By repeating the reasoning with $C_{1},C_{1}+g$ replaced by $C_{2}+g_{2},C_{2}+g_{2}+g$ we find $C_{3}$, a connected component of $E_{1}$, and $g_{3}\in G$ such that $C_{3}+g_{3}$ has a common vertex with $C_{2}+g_{2}$ and $C_{2}+g_{2}+g$. Moreover, $C_{3}+g_{3}+g$ is adjacent to $C_{3}+g_{3}$. Now let $u_{1},v,w,-u_{1},-v,-w$ be the vectors representing the ordered edges of $C_{1}$, let $u_{2},v,w,-u_{2},-v,-w$ be the ordered edges of $C_{2}$, and $u_{3},v,w,-u_{3},-v,-w$ be the ordered edges of $C_{3}$. Notice that $u_{1}=t_{1}u$, $u_{2}=t_{2}u$, $u_{3}=t_{3}u$ all have the same direction $u$. The hexagons $C_{1}+kg$, $C_{2}+g_{2}+kg$, $C_{3}+g_{3}+kg$ with $k\in\mathbb{Z}$ form three parallel strips with direction $g$ whose interfaces are zig-zags following the vectors $v$ and $w$. We can swap the second and third strip to obtain a new tiling with the same areas and with a whole interface removed. This is contrary to the minimality of $\mathbf{T}$} unless $C_{3}=C_{1}=E_{1}$. Analogously one finds $C_{2}=E_{2}$. Hence $E_{1}$ and $E_{2}$ are both connected. Uniqueness is straightforward, concluding the proof. ∎
222211222222
@@
Hence $C$ must have at least a vertex $v$ of type 1-2-2. The flat edge in this vertex has a corresponding opposite flat edge, obtained by means of a translation $g\in G$, which also must have a 1-2-2 corresponding vertex $w=v+g$. Following the curved edge of $C$ in $v$ we arrive at another vertex of type 1-2-2, then following the straight edge of $C$ we arrive at a vertex which can be either of type 1-2-2 or of type 2-2-2. In the first case we can follow the curved edge of $C$ to reach another vertex of type 1-2-2. Then, following the straight edge of $C$ we arrive in a vertex which is separated from $w$ by the curved edge of $w$.
@@
In both configurations the flat edges of $C$ can be prolonged to obtain a flat hexagon $P$ containing $C$ such that $\left\{P+g\colon g\in G\right\}$ is a tiling of $\mathbb{R}^{2}$. This means that $E_{2}=C$. Each vertex of $P$ is shared by $P$ and by $P+g$ and $P+h$ with $g,h\in G$. If the vertex of $P$ is not a vertex of $C$ we easily deduce that $P\setminus C$ is contained in a component of $E_{1}$ which is a Reauleaux triangle (i.e., a curvilinear triangle with three equal angles of 120 degrees).
@@
The configuration $(3,3;12)$ is a stationary configuration but we claim that it is not isoperimetric. In fact, it is possible to increase the size of the larger of the two Reauleaux triangle (or any of them if they are equal) and decrease the size of the other Reauleaux triangle so that the total area enclosed by the two triangles remains constant while the perimeter decreases. This is because the contribution to the perimeter of a Reauleaux triangle (the difference between the perimeter of the triangle and the length of the corresponding tripod) is increasing and proportional to the linear size of the triangle, while the enclosed area is proportional to the square of the linear size. So, if we have two Reauleaux triangles of respective sizes $r$ and $s$, the total volume is kept fixed if $k=r^{2}+s^{2}$ is constant while the contribution to the total perimeter is proportional to
@@
Notice that the configurations $(3;9)$ is, for a general lattice, not unique because we can choose to place the Reauleaux triangle in two different vertices of the hexagon. Notice that Theorem 2.4 combined with Theorem 2.6 directly implies our first main result Theorem $1.$